OpenStudy (thatgirl21):

Needing help with pre-calc for business study guide. My teacher doesn't explain it well so I need to know how to answer the questions, not just answers. 1. Evaluate each function at the given value: a.) if f(x)=11-7x, then find f(-3) b.) if g(x,y)=3x+ sqrt of(2y+1), then find g(6,12) c.) if h(x)=4^x + (x+1)/(x-1), then find h(3)

OpenStudy (thatgirl21):

for a I got f(-3)=-10, b I got g(6,12)=23, and c I got h(3)=65 2/3. is this right?

OpenStudy (loser66):

a,b ok c, why 2/3?

zepdrix (zepdrix):

Hmm a doesn't look quite right...\[\large\rm f(x)=11-7x\]\[\large\rm f(-3)=11-7(-3)\]That ends up being 11+21, yes? Double negative. Oh also, \(\large\rm \color{royalblue}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (loser66):

hehehe... zepdrix is soooo good.

OpenStudy (thatgirl21):

for c i got h(3)=64+5/3 and i simplified. not sure if that's even right

OpenStudy (thatgirl21):

oh I see what I did on a now. Forgot my negative sign and wrote it as a positive.

OpenStudy (loser66):